# Communication Constellation Design of Minimum Number of Satellites with Continuous Coverage and Inter-Satellite Link

### v1

@@ 5.2.2 Common Ground-track Constellation
The relative motion equations (48), (50), and (51) implies that the relative motion of two orbits with the same altitude and inclination is the function of the inclination, relative RAAN, and relative argument of latitude. Let us define the set of temporal location $\tau_{k}$ in Eq. (26) as $\boldsymbol{\tau}_{k}$, then $\boldsymbol{\tau}_{k,qs}$ and $\boldsymbol{\tau}_{k,bilp}$ are calculated as Eqs. (53) and (54). The temporal location $\tau_{k}$ derives the RAAN $\Omega_{k}$ as Eq. (55)[7].
@@
where the subscript ’0’ means the variable is relevant to the seed satellite. The equation (25) describes the relationship between $\Omega_{k}$ obtains $M_{k}$.
Let us define $\Delta\tau_{k}$, the subtraction of the consecutive $\tau_{k}$, as
@@
Thus, the equations (53), (54), and (56) calculates $\Delta\tau_{k}$ for the two CGT constellations as
@@
The equation (55) derives the relative RAAN $\Delta\Omega_{k}$ and relative mean anomaly $\Delta M_{k}$ as
@@
The minimum and maximum relative distances in Eqs. (46) and (47) are depicted in Figure 10. The left and right panels appear almost the same because the subtraction between the minimum and maximum distances are less than 1000 km. The minimum and maximum distances for the quasi-symmetric CGT constellation are 13854.32 and 13886.49 km, respectively. The BILP CGT constellation have a variety of $\Delta\tau_{k}$. When $\Delta\tau_{k}$ is 8, the minimum and maximum distances are 13164.56 and 13191.33 km.
@@ 6 Conclusion
This paper investigates the continuous coverage analysis methods and inter-satellite link connectivity analysis methods for the communication satellite constellation. The bounded Voronoi diagram designs the homogeneous constellation to have continuous coverage for regional or global coverage. The APC decomposition can be implemented for the CGT constellation’s coverage analysis. The relative motion in the adjacent orbital planes derives the analytic solutions that should be constrained within the inter-satellite link range. The Walker-Delta constellation and two types of CGT constellations are applied as examples. As a result, Walker-Delta constellation design method consists of more satellites than CGT constellations for single-fold coverage. However, the relative motion range is shorter and more consistent, which implies that the Walker-Delta constellation has advantages for inter-satellite links. The quasi-symmetric and BILP methods are suggested for the CGT constellation design method. The BILP optimal CGT constellation has an asymmetric configuration but achieves less number of satellites. The relative motion for BILP constellation have a variety of ranges due to its asymmetricity, but satellites are located closer than the quasi-symmetric constellation. In summary, the BILP constellation is advantageous for the required number of satellites for single-fold coverage. However, the Walker-Delta constellation may have a shorter and more rigorous relative motion range that is beneficial for inter-satellite links.
@@ 7 Acknowledgment
Any acknowledgments by the author may appear here. The acknowledgments section is optional. This work was supported by Korea Research Institute for defense Technology planning and advancement (KRIT) grant funded by the Korea government (DAPA (Defense Acquisition Program Administration)) (KRIT-CT-22-040, Heterogeneous Satellite constellation based ISR Research Center, 2024).

### v2

@@ 5.2.2 Common Ground-track Constellation
The relative motion equations (48), (50), and (51) imply that the relative motion of two orbits with the same altitude and inclination is a function of the inclination, relative RAAN, and relative argument of latitude. Let us define a set of temporal location $\tau_{k}$ in Eq. (26) as $\boldsymbol{\tau}_{k}$. Then, $\boldsymbol{\tau}_{k,qs}$ and $\boldsymbol{\tau}_{k,bilp}$ are calculated as Eqs. (53) and (54). The temporal location $\tau_{k}$ derives the RAAN $\Omega_{k}$ as shown in Eq. (55)[7].
@@
where the subscript ‘0’ means that the variable is relevant to the seed satellite. The equation (25) describes the relationship between $\Omega_{k}$ and $M_{k}$.
Let us define $\Delta\tau_{k}$ as the difference between consecutive $\tau_{k}$ values:
@@
Thus, the equations (53), (54), and (56) are used to calculate $\Delta\tau_{k}$ for the two CGT constellations as
@@
The equation (55) derives the relative RAAN $\Delta\Omega_{k}$ and the relative mean anomaly $\Delta M_{k}$ as
@@
The minimum and maximum relative distances in Eqs. (46) and (47) are depicted in Figure 10. The left and right panels appear almost identical because the difference between the minimum and maximum distances is less than 1000 km. The minimum and maximum distances for the quasi-symmetric CGT constellation are 13854.32 and 13886.49 km, respectively. The BILP CGT constellation has various values of $\Delta\tau_{k}$. When $\Delta\tau_{k}$ is 8, the minimum and maximum distances are 13164.56 and 13191.33 km.
@@ 6 Conclusion
This paper investigates the continuous coverage analysis methods and inter-satellite link connectivity analysis methods for communication satellite constellations. The bounded Voronoi diagram designs a homogeneous constellation to ensure continuous regional and global coverage. The APC decomposition can be implemented for the CGT constellation’s coverage analysis. The relative motion in adjacent orbital planes derives the analytical solutions that should be constrained within the inter-satellite link range. The Walker-Delta constellation and two types of CGT constellations are applied as examples. As a result, the Walker-Delta constellation design method consists of more satellites than CGT constellations for a single-fold coverage. However, the relative motion range is shorter and more consistent, which implies that the Walker-Delta constellation has advantages for inter-satellite links. The quasi-symmetric and BILP methods have been suggested for the CGT constellation design method. The BILP optimal CGT constellation has an asymmetric configuration but achieves a smaller number of satellites. The relative motion of BILP constellation has a variety of ranges due to its asymmetricity, but satellites are located closer than the quasi-symmetric constellation. In summary, the BILP constellation is advantageous in terms of the number of satellites required for a single-fold coverage. However, the Walker-Delta constellation may have a shorter and more rigorous relative motion range, which is beneficial for inter-satellite links.
@@ 7 Acknowledgment
This work was supported by the Korea Research Institute for defense Technology Planning and Advancement (KRIT) grant funded by the Korean government (DAPA (Defense Acquisition Program Administration)) (KRIT-CT-22-040, Heterogeneous Satellite Constellation-based ISR Research Center, 2024).